OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? x+3y=12 -4x+3y=-3 A. (0, 4) B. (6, 2) C. (3, 3) D. (0, −1)

OpenStudy (anonymous):

C

OpenStudy (anonymous):

-x-3y=-12 -4x+3y=-3 ----------- -5x=-15 x=3

OpenStudy (anonymous):

3+3y=12 3y=9 y=3

OpenStudy (anonymous):

thanks so much!!